What would be the theoretical yield of this Grignard reaction with the product being benzoic acid?

0.70 mL of bromobenzene

0.15 g Mg

5g dry ice

crude benzoic acid recovered: 0.08 g
Answer #1

Solution:

Formation of benzoic acid from bromobenzene via grignard reagent is given as,

C6H5Br + Mg = C6H5MgBr

C6H5MgBr + CO2 + H2O = C6H5COOH + Mg(OH)Br

Calculation of theoretical yield :

Density of bromobenzene = 1.5 g /mL

Hence, mass = volume x density

= 0.70 mL x 1.5 g mL^-1 = 1.05 g

Number of moles of Bromobenzene = mass./ molar mass

= 1.05 g / 157.02 g mol-1 = 0.0067 mol

Number of moles of Mg = mass./ molar mass

= 0.15 g / 24.305 g mol-1 = 0.0062 mol

Number of moles of dry ice = mass./ molar mass

= 5.0 g / 18 g mol-1 = 0.278 mol

From the above given equations, it can be seen that all the reactants reacted in 1:1 molar ratio and minimum number of moles are obtained for Mg, hence Mg will be limiting reagent.

Hence, number of moles of Mg will be equal to the number ilof moles of Benzoic acid

Thus, number of mol of benzoic acid = 0.0062 mol

Hence, theoretical yield = number of mol x molar mass

= 0.0062 mol x 122.12 g mol-1 = 0.757 g